# Definability of mad families of vector spaces and two local Ramsey theories

### v4

@@ 4.2. Ideals and semicoideals generated by almost disjoint families
    $\langle B_{n}\rangle\subseteq\langle Y\rangle$.
Note that these two properties imply that $\langle Y\rangle\in\mathcal{I}^{++}(\mathcal{A})$.
Proof.
Since $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$, $\langle Y\rangle$ is big, so there exists some $B_{0}^{\prime}$ such that $\langle B_{0}^{\prime}\rangle\subseteq\langle Y\rangle$. By our assumption, $B_{0}^{\prime}$ is compatible with some $A_{0}\in\mathcal{A}$. Let $B_{0}$ be such that $\langle B_{0}\rangle\subseteq\langle B_{0}^{\prime}\rangle\cap\langle A_{0}\rangle$. Since $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$, $\langle Y\rangle\setminus\langle A_{0}\rangle$ is big. Given an ordinal $\alpha$, suppose that we have defined an infinite disjoint family $\{B_{\beta}:\beta<\alpha\}$ and $\{A_{\beta}:\beta<\alpha\}\subseteq\mathcal{A}$ such that $B_{\beta}\leq A_{\beta}$ for all $\beta<\alpha$, and $\langle B_{\beta}\rangle\subseteq\langle Y\rangle$ for all $\beta<\alpha$. If $Y\setminus\bigcup_{\beta<\alpha}\langle B_{\beta}\rangle$ is small, we terminate the construction. Otherwise, we let $B_{\alpha}^{\prime}$ be such that $\langle B_{\alpha}^{\prime}\rangle\subseteq Y\setminus\bigcup_{\beta<\alpha}\langle B_{\beta}\rangle$, and let $A_{\alpha}\in\mathcal{A}$ be compatible with $B_{\alpha}^{\prime}$. We then let $B_{\alpha}$ be such that $\langle B_{\alpha}\rangle\subseteq\langle B_{\alpha}^{\prime}\rangle\cap\langle A_{\alpha}\rangle$. Since $\langle Y\rangle$ is countable and the $B_{\alpha}$’s are disjoint, this induction must terminate at some countable stage $\alpha$, and since $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$, $\{A_{\beta}:\beta<\alpha\}$ must be infinite. ∎
4.3. Selectivity properties revisited
We now return to studying the combinatorial properties of the semicoideals $\mathcal{H}^{-}(\mathcal{A})$ and $\mathcal{H}(\mathcal{A})$. In Mathias’ process of proving that there are no analytic mad families of $\mathbb{N}$$^{\infty}$, he proved that if $\mathcal{A}$ is an almost disjoint family in $\mathbb{N}$$^{\infty}$, then $\mathcal{H}^{-}(\mathcal{A})$ is a selective coideal. In [S19], Smythe proved that if $\mathcal{A}$ is a mad family in $(E^{[\infty]},\leq,r)$, then $\mathcal{H}(\mathcal{A})$ is a (p)-semicoideal. We prove a few extensions of this result.
Proposition 4.12.
Let $\mathcal{A}\subseteq E^{[\infty]}$ be an almost disjoint family. Then $\mathcal{H}^{-}(\mathcal{A})$ is a (p)-semicoideal.
Proof.
Let $(B_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}^{-}(\mathcal{A})$. If there exists some $C$ which weakly diagonalises $(B_{n})_{n<\omega}$ and $C$ is almost disjoint from every element of $\mathcal{A}$, then $C\in\mathcal{H}^{-}(\mathcal{A})$ and we are done, so assume otherwise. Let $C_{0}$ be a weak diagonalisation of $(B_{n})_{n<\omega}$. By our assumption, $C_{0}$ is compatible with some $A_{0}\in\mathcal{A}$, so by taking a further block subspace of $C_{0}$ if necessary, we may assume that $C_{0}\leq A_{0}$. Now assume, for the induction hypothesis, that we have defined $C_{k},A_{k}$ for $k<m$ such that:
@@ 4.3. Selectivity properties revisited
    $B_{n}^{m}$ is almost disjoint from $A_{k}$ for all $k<m$.
@@
For any $n$, since $B_{n}\in\mathcal{H}^{-}(\mathcal{A})$ there exists some $D_{n}\leq B_{n}$ that is disjoint from $A_{k}$ for all $k<m$. Let $\Gamma:\omega\to\omega$ be any infinite-to-one surjection. We define $D=(x_{n})_{n<\omega}$ as follows: Assuming that $(x_{i})_{i<n}$ has been defined so far, by Lemma 4.3 we may let $N$ be large enough so that for all $k<m$ and $x\in\langle B_{0}/N\rangle$ such that $x\notin\langle A_{k}\rangle$, we have that:
$\langle x_{0},\dots,x_{n-1},x\rangle\cap\langle A_{k}\rangle=\langle x_{0},\dots,x_{n-1}\rangle\cap\langle A_{k}\rangle$
We then let $x_{n}\in\langle D_{\Gamma(n)}\rangle\cap\langle B_{0}/N\rangle$ be any element. $D$ is almost disjoint from $A_{k}$ for all $k<m$, as:
$\langle D\rangle\cap\langle A_{k}\rangle$ $=\bigcup_{i<\omega}\langle x_{0},\dots,x_{i}\rangle\cap\langle A_{k}\rangle=\langle x_{0},\dots,x_{k-1}\rangle\cap\langle A_{k}\rangle.$
Let $B_{n}^{m}:=(x_{i})_{i<\omega,\Gamma(i)\geq n}$ Then $B_{n}^{m}\leq B_{n}$, $(B_{n}^{m})_{n<\omega}$ is a $\leq$-decreasing sequence, and $B_{n}^{m}\leq D$, so $B_{n}^{m}$ is almost disjoint from $A_{k}$ for all $k\leq m$. ∎
Let $C_{m}\leq B_{m}^{m}$ be a weak diagonalisation of $(B_{n}^{m})_{n<\omega}$. By our assumption at the beginning, there exists some $A_{m}\in\mathcal{A}$ which is compatible with $C_{m}$, and we may assume that $C_{m}\leq A_{m}$. Since $C_{m}\leq B_{m}^{m}$, and $B_{m}^{m}$ is almost disjoint from $A_{k}$ for all $k<m$, we have that $A_{k}\neq A_{m}$ for all $k<m$. Having completed the induction, we let $C$ be a weak diagonalisation of $(C_{m})_{m<\omega}$. Then $C\leq^{*}C_{0}\leq^{*}B_{n}$ for any $n$, so $C$ weak diagonalises $(B_{n})_{n<\omega}$. Since $C\leq^{*}C_{m}\leq A_{m}$ for any $m$, $C$ is compatible with infinitely many elements of $\mathcal{A}$, so $C\in\mathcal{H}^{-}(\mathcal{A})$. ∎
@@
Let $\mathcal{A}\subseteq E^{[\infty]}$ be an almost disjoint family. Then $\mathcal{H}^{-}(\mathcal{A})$ is a selective semicoideal.
@@
Suppose there is no $B\leq A$ which is almost disjoint from every element of $\mathcal{A}$. By Lemma 4.11, there exists an infinite $\{A_{n}:n<\omega\}\subseteq\mathcal{A}$ and an infinite disjoint family $\{B_{n}:n<\omega\}$ such that $\langle B_{n}\rangle\subseteq\langle A\rangle\cap\langle A_{n}\rangle$ for all $n$. Let $\Gamma:\omega\to\omega$ be any infinite-to-one surjection. We write $A=(x_{n})_{n<\omega}$, and define $B=(y_{n})_{n<\omega}$ as follows: If $(y_{k})_{k<n}$ has been defined, we let $m$ be the least integer such that $\langle r_{n}(B)\rangle\subseteq\langle x_{0},\dots,x_{\min(I_{m})-1}\rangle$. We let $y_{n}\in\langle B_{\Gamma(n)}/\max(I_{m})\rangle$ be any element. Then $(y_{n})_{n\in\Gamma^{-1}(m)}$ witnesses that $B$ and $B_{m}$ are compatible for all $m<\omega$, and since $y_{n}\in\langle A\rangle$ for all $n$, $B\leq A$ and $B\in\mathcal{H}^{-}(\mathcal{A})$. ∎
5. Construction of mad families of subspaces
@@
A mad family $\mathcal{A}\subseteq E^{[\infty]}$ is full if $\mathcal{H}(\mathcal{A})$ is full.
@@ 5.1. Block-splitting families
$\mathfrak{b}:=\{|\mathcal{F}|:\mathcal{F}\subseteq\omega^{\omega}\text{ is unbounded}\}.$
@@
$\mathfrak{s}:=\{|\mathcal{S}|:\mathcal{S}\subseteq[\omega]^{\omega}\text{ is a splitting family}\}.$
It turns out that both of these cardinals are closely related to block-splitting families. If $\mathcal{P}=\{P_{n}:n<\omega\}$ is an interval partition, we say that a set $S\in[\omega]^{\omega}$ splits $\mathcal{P}$ if $\{n<\omega:P_{n}\subseteq S\}$ and $\{n<\omega:P_{n}\cap S=\emptyset\}$ are both infinite. A block-splitting family is a family $\mathcal{S}\subseteq[\omega]^{\omega}$ such that every interval partition is block-split by some $S\in\mathcal{S}$. The following lemma is due to Kamburelis-Weglorz.
Lemma 5.2 ([KW96]).
@@
We abbreviate $\mathfrak{bs}=\max\{\mathfrak{b},\mathfrak{s}\}$. By Corollary 2.11 of [S19], the identity $\operatorname{non}(\mathcal{M})\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ holds for any countable field $\mathbb{F}$. Since $\mathfrak{b}\leq\operatorname{non}(\mathcal{M})$ and $\mathfrak{s}\leq\operatorname{non}(\mathcal{M})$ (see, for instance, [handbook]), we may conclude that:
@@
Let $A,B\in E^{[\infty]}$.
  1. (1)
    If $\langle A\rangle\cap\langle B\rangle$ is small, then $\langle A\rangle\cap\langle B\rangle$ is very small.
  2. (2)
    If $\langle A\rangle\setminus\langle B\rangle$ is small, then $\langle A\rangle\setminus\langle B\rangle$ is very small.
Consequently, it is not possible to have both $\langle A\rangle\cap\langle B\rangle$ and $\langle A\rangle\setminus\langle B\rangle$ small.
Proof.
Assume that $\langle A\rangle\cap\langle B\rangle$ is small. Let $Y\subseteq E$, and suppose that $\langle C\rangle\subseteq(\langle A\rangle\cap\langle B\rangle)\cup Y$ for some $C=(x_{n})_{n<\omega}\in E^{[\infty]}$. Since $\langle A\rangle\cap\langle B\rangle$ is a finite-dimensional space, we have that $M:=\max(\operatorname{supp}[\langle A\rangle\cap\langle B\rangle])<\infty$. Let $N$ be large enough so that $\operatorname{supp}(x_{n})>M$. Then $\langle C/N\rangle\cap(\langle A\rangle\cap\langle B\rangle)=\emptyset$, so $\langle C/N\rangle\subseteq Y$. Therefore, $Y$ is big.
Now assume that $\langle A\rangle\setminus\langle B\rangle$ is small. Let $Y\subseteq E$, and suppose that $\langle C\rangle\subseteq(\langle A\rangle\setminus\langle B\rangle)\cup Y$. We consider three cases. For the first case, suppose that $C$ is almost disjoint from $A$. By Lemma 4.9, let $N$ be large enough so that $\langle C/N\rangle\cap\langle A\rangle=\emptyset$. Then this implies that $\langle C/N\rangle\subseteq Y$, so $Y$ is big.
For the second case, suppose that $C$ is compatible with $B$. Let $D\in E^{[\infty]}$ be such that $\langle D\rangle\subseteq\langle C\rangle\cap\langle B\rangle$. Then $\langle D\rangle\subseteq(\langle A\rangle\setminus\langle B\rangle)\cup Y$, and since $\langle D\rangle\cap(\langle A\rangle\setminus\langle B\rangle)=\emptyset$, we have that $\langle D\rangle\subseteq Y$. Therefore, $Y$ is big.
For the last case, suppose that $C$ is compatible with $A$ but almost disjoint from $B$. Let $D\in E^{[\infty]}$ be such that $\langle D\rangle\subseteq\langle C\rangle\cap\langle A\rangle$, and note that $D$ is also almost disjoint from $B$. By Lemma 4.9, let $N$ be large enough so that $\langle D/N\rangle\cap\langle B\rangle=\emptyset$. Since $D/N\leq A$, this implies that $\langle D/N\rangle\subseteq\langle A\rangle\setminus\langle B\rangle$, contradicting our assumption that $\langle A\rangle\setminus\langle B\rangle$ is small. ∎
Lemma 5.5.
There exists a collection $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ of subsets of $E$ such that for all almost disjoint families $\mathcal{A}\subseteq E^{[\infty]}$ and $B\in\mathcal{H}^{-}(\mathcal{A})$, the following properties hold:
@@
Property (i): Suppose that for all $\leq B, CC$ is compatible with some element of $\mathcal{A}$. Suppose that $\langle B\rangle\cap Z_{\alpha}\in\mathcal{I}(\mathcal{A})$ and $\langle B\rangle\cap Z_{\alpha}^{c}\in\mathcal{I}(\mathcal{A})$. Let $\{A_{i}:i<n\}\subseteq\mathcal{A}$ such that $(\langle B\rangle\cap Z_{\alpha})\setminus\bigcup_{i<n}\langle A_{i}\rangle$ and $(\langle B\rangle\cap Z_{\alpha}^{c})\setminus\bigcup_{i<n}\langle A_{i}\rangle$ are small. Since $B\in\mathcal{H}^{-}(\mathcal{A})$, there exists some $C\leq B$ such that:
$\langle C\rangle\subseteq\langle B\rangle\setminus\bigcup_{i<n}\langle A_{i}\rangle.$
@@
$\langle C\rangle\cap Z_{\alpha}\subseteq(\langle B\rangle\cap Z_{\alpha})\setminus\bigcup_{i<n}\langle A_{i}\rangle,$
$\langle C\rangle\cap Z_{\alpha}^{c}\subseteq(\langle B\rangle\cap Z_{\alpha}^{c})\setminus\bigcup_{i<n}\langle A_{i}\rangle,$
@@
Property (ii): This follows from Lemma 2.3, where we may let $C_{0}^{\alpha}\in E^{[\infty]}$ be such that $\langle C_{0}^{\alpha}\rangle=\langle B\rangle\cap\langle(e_{n})_{n\in S_{\alpha}}\rangle$.
Property (iii): Suppose that $\langle B\rangle\cap Z_{\alpha}^{c}\in\mathcal{I}^{+}(\mathcal{A})$. If there exists some $C\in\mathcal{R}$ such that $\langle C\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$ which is almost disjoint from every element of $\mathcal{A}$, then $C\in\mathcal{H}^{-}(\mathcal{A})$ and we are done, so assume otherwise. In a way similar to the one presented in the proof of Proposition 4.12, we define $A_{n},B_{n}$ for $n<\omega$ such that $\bigcup_{n<\omega}\langle B_{n}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$, $\{B_{n}:n<\omega\}$ is a disjoint family, and $A_{n}$ is compatible with $B_{n}$ for all $n$. Let $C_{n}$ be such that $\langle C_{n}\rangle\subseteq\langle A_{n}\rangle\cap\langle B_{n}\rangle$.
For each $n$, we write $C_{n}:=(x_{m}^{n})_{m<\omega}$. Since $\langle C_{n}\rangle\subseteq Z_{\alpha}^{c}$, this means that for all $m$, there exists some $l_{m,n}\notin S_{\alpha}$ and some $y_{m}^{n},z_{m}^{n}\in\langle C_{n}\rangle\cup\{0\}$ such that $y_{m}^{n}<e_{l_{m,n}}<z_{m}^{n}$, and $x_{m}^{n}=y_{m}^{n}+\lambda_{m,n}e_{l_{m,n}}+z_{m}^{n}$ for some $\lambda_{m,n}\neq 0$. Now $\Gamma:\omega\to\omega\times\omega$ be any function such that:
@@
    if $k<k^{\prime}$, $\Gamma(k)=(n,m)$ and $\Gamma(k^{\prime})=(n^{\prime},m^{\prime})$, then $x_{n}^{m}<x_{n^{\prime}}^{m^{\prime}}$.
This implies that $C:=\left(x_{\Gamma(k)(1)}^{\Gamma(k)(0)}\right)_{k<\omega}$ is a well-defined infinite block sequence, where $\Gamma(k)=(\Gamma(k)(0),\Gamma(k)(1))$ such that $C$ is compatible with every $C_{n}$, so $C\in\mathcal{H}^{-}(\mathcal{A})$. Furthermore, we have that:
@@
so $C\leq B$. Finally, if $x\in\langle C\rangle$, then $x=y+\lambda e_{l}+z$ for some $\lambda\neq 0$, so $x\notin Z_{\alpha}$. In other words, $\langle C\rangle\subseteq Z_{\alpha}^{c}$. Therefore, $C\in\mathcal{H}^{-}(\mathcal{A})$ and $\langle C\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$, as desired.
@@
$P_{0}$ $:=\{0,\dots,\max(\operatorname{supp}(x_{n}))\},$
$P_{n+1}$ $:=\{\max(\operatorname{supp}(x_{n}))+1,\dots,\max(\operatorname{supp}(x_{n+1})\}.$
@@
We have that $\mathcal{P}:=\{P_{n}:n<\omega\}$ is an interval partition such that $\operatorname{supp}(x_{n})\subseteq P_{n}$ for all $n$. Since $\{S_{\alpha}:\alpha<\mathfrak{bs}\}$ is a block-splitting family, there exists some $\alpha<\mathfrak{bs}$ that splits $\mathcal{P}$. Let $T_{0}:=\{n<\omega:P_{n}\subseteq S_{\alpha}\}$ and $T_{1}:=\{n<\omega:P_{n}\cap S_{\alpha}^{c}\}$. Define $B_{0}:=(x_{n})_{n\in T_{0}}$ and $B_{1}:=(x_{n})_{n\in T_{1}}$. Since $B_{0},B_{1}\leq B$, and $B$ is almost disjoint from every element of $\mathcal{A}$, so are $B_{0},B_{1}$ and so $B_{0},B_{1}\in\mathcal{H}^{-}(\mathcal{A})$. If $n\in T_{0}$, then $\operatorname{supp}(x_{n})\subseteq P_{n}\subseteq S_{\alpha}$, $x_{n}\in(e_{n})_{n\in S_{\alpha}}$. Therefore, $\langle B_{0}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}$. By a similar reasoning, we may also conclude that $\langle B_{1}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$.

### v5

@@ 4.3. Selectivity properties revisited
    $\langle C_{k}\rangle\subseteq V_{k}$ and $V_{k}\in\mathcal{A}$.
@@
    $B_{n}^{m}$ is almost disjoint from $V_{k}$ for all $k<m$.
@@
For any $n$, since $B_{n}\in\mathcal{H}^{-}(\mathcal{A})$ there exists some $D_{n}\leq B_{n}$ that is disjoint from $V_{k}$ for all $k<m$. Let $\Gamma:\omega\to\omega$ be any infinite-to-one surjection. We define $D=(x_{n})_{n<\omega}$ as follows: Since $D_{\Gamma(0)}$ is almost disjoint from $V_{k}$ for all $k<m$, we may let $x_{0}\in\langle D_{\Gamma(0)}\rangle\setminus\bigcup_{k<m}V_{k}$ be any vector (which exists by Lemma 4.8). This implies that $\langle x_{0}\rangle\cap V_{k}=\emptyset$ for all $k<m$. Now assume that $(x_{i})_{i<n}$ has been defined, and for all $k<m$:
$\langle x_{0},\dots,x_{n-1}\rangle\cap V_{k}=\emptyset.$
By Claim 5 in the proof of Lemma 4.4, for each $k<m$ since $D_{\Gamma(n)}$ and $V_{k}$ are almost disjoint, we have that $D_{\Gamma(n)}$ and $V_{k}+\langle x_{0},\dots,x_{n-1}\rangle$ are also almost disjoint. Let $x_{n}\in\langle D_{\Gamma(n)}\rangle\setminus\bigcup_{k<m}(V_{k}+\langle x_{0},\dots,x_{n-1}\rangle)$ be any vector (again, which exists by Lemma 4.8). We claim that for all $k<m$:
$\langle x_{0},\dots,x_{n}\rangle\cap V_{k}=\emptyset.$
Otherwise, for any $k<m$ let $y\in\langle x_{0},\dots,x_{n}\rangle\cap V_{k}$ be any vector, so $y=\sum_{i\leq n}\xi_{i}x_{i}$. Since $\langle x_{0},\dots,x_{n-1}\rangle\cap V_{k}=\emptyset$, we have that $\xi_{n}\neq 0$. But this implies that:
$x_{n}=\frac{1}{\xi_{n}}\left(y-\sum_{i<n}\xi_{i}x_{i}\right)\in V_{k}+\langle x_{0},\dots,x_{n-1}\rangle,$
contradicting our choice of $x_{n}$. Then $D=(x_{n})_{n<\omega}$ is disjoint from $V_{k}$ for all $k<m$, as:
$\langle D\rangle\cap V_{k}$ $=\bigcup_{i<\omega}\langle x_{0},\dots,x_{i}\rangle\cap V_{k}=\emptyset.$
Let $B_{n}^{m}:=(x_{i})_{\Gamma(i)\geq n}$. Since $\Gamma$ is infinite-to-one, $B_{n}^{m}$ is an infinite block sequence for all $n,m$. Then $B_{n}^{m}\leq B_{n}$, $(B_{n}^{m})_{n<\omega}$ is a $\leq$-decreasing sequence, and $B_{n}^{m}\leq D$, so $B_{n}^{m}$ is almost disjoint from $V_{k}$ for all $k<m$. ∎
Let $C_{m}\leq B_{m}^{m}$ be a weak diagonalisation of $(B_{n}^{m})_{n<\omega}$. By our assumption at the beginning, there exists some $V_{m}\in\mathcal{A}$ which is compatible with $C_{m}$, and we may assume that $\langle C_{m}\rangle\subseteq V_{m}$. Since $C_{m}\leq B_{m}^{m}$, and $B_{m}^{m}$ is almost disjoint from $V_{k}$ for all $k<m$, we have that $V_{k}\neq V_{m}$ for all $k<m$. This completes the inductive definition of $C_{m},V_{m}$ for $m<\omega$.
Let $\Gamma:\omega\to\omega$ be an infinite-to-one surjection, and define $C:=(y_{i})_{i<\omega}$ as follows: Let $y_{0}\in\langle C_{\Gamma(0)}\rangle$, and let $y_{i}\in\langle C_{\Gamma(i)}/i\rangle$ be any vector such that $y_{i-1}<y_{i}$. Observe that for all $m$, we have that:
$\langle y_{i}:\Gamma(i)=m\rangle\subseteq\langle C\rangle\cap\langle C_{m}\rangle\subseteq\langle C\rangle\cap V_{m},$
so $C$ is compatible with $V_{m}$ for all $m$. Thus, $C\in\mathcal{H}^{-}(\mathcal{A})$. It remains to show that $C$ weakly diagonalises $(B_{n})_{n<\omega}$. Fix any $n$, and let $N$ be large enough so that $C_{k}/N\leq B_{n}$ for all $k<n$. For all $i\geq N$, we have that $y_{i}\in\langle C_{\Gamma(i)}/i\rangle\subseteq\langle C_{\Gamma(i)}/N\rangle$. If $\Gamma(i)<n$, then $y_{i}\in\langle C_{\Gamma(i)}/N\rangle\subseteq\langle B_{n}\rangle$ by our choice of $N$. If $\Gamma(i)\geq n$, then $C_{\Gamma(i)}\leq B_{\Gamma(i)}^{\Gamma(i)}\leq B_{\Gamma(i)}\leq B_{n}$, so $y_{i}\in\langle B_{n}\rangle$. Therefore, $C/N\leq B_{n}$. ∎
@@
Let $\mathcal{A}$ be an infinite almost disjoint family of subspaces. Then $\mathcal{H}^{-}(\mathcal{A})$ is a selective semicoideal.
@@
Suppose there is no $B\leq A$ which is almost disjoint from every element of $\mathcal{A}$. By Lemma 4.13, there exists an infinite $\{V_{n}:n<\omega\}\subseteq\mathcal{A}$ and an infinite disjoint family $\{B_{n}:n<\omega\}$ such that $\langle B_{n}\rangle\subseteq\langle A\rangle\cap V_{n}$ for all $n$. Let $\Gamma:\omega\to\omega$ be any infinite-to-one surjection. We write $A=(x_{n})_{n<\omega}$, and define $B=(y_{n})_{n<\omega}$ below $A$ as follows: If $(y_{k})_{k<n}$ has been defined, we let $m$ be the least integer such that $\langle y_{k}:k<n\rangle\subseteq\langle x_{0},\dots,x_{\min(I_{m})-1}\rangle$. We let $y_{n}\in\langle B_{\Gamma(n)}/\max(I_{m})\rangle$ be any element. Then $(y_{n})_{n\in\Gamma^{-1}(m)}$ witnesses that $B$ and $B_{m}$ are compatible for all $m<\omega$, and since $y_{n}\in\langle A\rangle$ for all $n$, $B\leq A$ and $B\in\mathcal{H}^{-}(\mathcal{A})$. ∎
5. Constructions of mad families of subspaces
@@
A mad family $\mathcal{A}$ of subspaces is full if $\mathcal{H}(\mathcal{A})$ is full (see Definitions 3.3 and 4.9).
@@ 5.1. Block-splitting families
$\mathfrak{b}:=\min\{|\mathcal{F}|:\mathcal{F}\subseteq\omega^{\omega}\text{ is unbounded}\}.$
@@
$\mathfrak{s}:=\min\{|\mathcal{S}|:\mathcal{S}\subseteq[\omega]^{\omega}\text{ is a splitting family}\}.$
It turns out that both of these cardinals are closely related to block-splitting families. If $\mathcal{P}=\{P_{n}:n<\omega\}$ is an interval partition, we say that a set $S\in[\omega]^{\omega}$ splits $\mathcal{P}$ if $\{n<\omega:P_{n}\subseteq S\}$ and $\{n<\omega:P_{n}\cap S=\emptyset\}$ are both infinite. A block-splitting family is a family $\mathcal{S}\subseteq[\omega]^{\omega}$ such that every interval partition is split by some $S\in\mathcal{S}$. The following lemma is due to Kamburelis-Weglorz.
Lemma 5.2 ([11]).
@@
We abbreviate $\mathfrak{bs}=\max\{\mathfrak{b},\mathfrak{s}\}$. By Corollary 2.11 of [24], the identity $\operatorname{non}(\mathcal{M})\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ holds for any countable field $\mathbb{F}$. Since $\mathfrak{b}\leq\operatorname{non}(\mathcal{M})$ and $\mathfrak{s}\leq\operatorname{non}(\mathcal{M})$ (see, for instance, [5]), we may conclude that:
@@
There exists a collection $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ of block subspaces of $E$ such that for all almost disjoint families $\mathcal{A}$ of subspaces and $B\in\mathcal{H}^{-}(\mathcal{A})$, the following properties hold:
@@
Property (i): Suppose that $\langle B\rangle\cap Z_{\alpha}\in\mathcal{I}(\mathcal{A})$ and $\langle B\rangle\cap Z_{\alpha}^{c}\in\mathcal{I}(\mathcal{A})$. Let $\{V_{i}:i<n\}\subseteq\mathcal{A}$ such that $(\langle B\rangle\cap Z_{\alpha})\setminus\bigcup_{i<n}V_{i}$ and $(\langle B\rangle\cap Z_{\alpha}^{c})\setminus\bigcup_{i<n}V_{i}$ are small. Since $B\in\mathcal{H}^{-}(\mathcal{A})$, there exists some $C\leq B$ such that:
$\langle C\rangle\subseteq\langle B\rangle\setminus\bigcup_{i<n}V_{i}.$
@@
$\langle C\rangle\cap Z_{\alpha}\subseteq(\langle B\rangle\cap Z_{\alpha})\setminus\bigcup_{i<n}V_{i},$
$\langle C\rangle\cap Z_{\alpha}^{c}\subseteq(\langle B\rangle\cap Z_{\alpha}^{c})\setminus\bigcup_{i<n}V_{i},$
@@
Property (ii): This follows from Lemma 2.3, where we may let $C_{\alpha}^{0}\in E^{[\infty]}$ be such that $\langle C_{\alpha}^{0}\rangle=\langle B\rangle\cap\langle(e_{n})_{n\in S_{\alpha}}\rangle$.
Property (iii): Suppose that $\langle B\rangle\cap Z_{\alpha}^{c}\in\mathcal{I}^{+}(\mathcal{A})$. If there exists some $C\in E^{[\infty]}$ such that $\langle C\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$ which is almost disjoint from every element of $\mathcal{A}$, then $C\in\mathcal{H}^{-}(\mathcal{A})$ and we are done, so assume otherwise. In a way similar to the one presented in the proof of Lemma 4.13, we define $V_{n},B_{n}$ for $n<\omega$ such that $\bigcup_{n<\omega}\langle B_{n}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$, $\{B_{n}:n<\omega\}$ is a disjoint family of block subspaces, and $V_{n}\in\mathcal{A}$ is compatible with $B_{n}$ for all $n$. Let $C_{n}$ be such that $\langle C_{n}\rangle\subseteq V_{n}\cap\langle B_{n}\rangle$.
For each $n$, we write $C_{n}:=(x_{n,m})_{m<\omega}$. Since $\langle C_{n}\rangle\subseteq Z_{\alpha}^{c}$, this means that for all $m$, $\lambda_{l_{n,m}}(x_{n,m})\neq 0$ for some $l_{m,n}\notin S_{\alpha}$. Now let $\Gamma:\omega\to\omega\times\omega$ be any function such that:
@@
    if $k<k^{\prime}$, $\Gamma(k)=(n,m)$ and $\Gamma(k^{\prime})=(n^{\prime},m^{\prime})$, then $x_{n,m}<x_{n^{\prime},m^{\prime}}$.
This implies that $C:=(x_{\Gamma(k)})_{k<\omega}$ is a well-defined infinite block sequences, where if $\Gamma(k)=(n,m)$ then $x_{\Gamma(k)}:=x_{n,m}$. Then $C$ is compatible with every $C_{n}$, so $C\in\mathcal{H}^{-}(\mathcal{A})$. Furthermore, we have that:
@@
so $C\leq B$. Finally, if $x\in\langle C\rangle$, then $\lambda_{l}(x)\neq 0$ for some $l\notin S_{\alpha}$, so $x\notin Z_{\alpha}$. In other words, $\langle C\rangle\subseteq Z_{\alpha}^{c}$. Therefore, $C\in\mathcal{H}^{-}(\mathcal{A})$ and $\langle C\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$, as desired.
@@
$P_{0}$ $:=\{0,\dots,\max(\operatorname{supp}(x_{0}))\},$
$P_{n+1}$ $:=\{\max(\operatorname{supp}(x_{n}))+1,\dots,\max(\operatorname{supp}(x_{n+1}))\}.$
@@
We have that $\mathcal{P}:=\{P_{n}:n<\omega\}$ is an interval partition such that $\operatorname{supp}(x_{n})\subseteq P_{n}$ for all $n$. Since $\{S_{\alpha}:\alpha<\mathfrak{bs}\}$ is a block-splitting family, there exists some $\alpha<\mathfrak{bs}$ that splits $\mathcal{P}$. Let $T_{0}:=\{n<\omega:P_{n}\subseteq S_{\alpha}\}$ and $T_{1}:=\{n<\omega:P_{n}\cap S_{\alpha}=\emptyset\}$. Define $B_{0}:=(x_{n})_{n\in T_{0}}$ and $B_{1}:=(x_{n})_{n\in T_{1}}$. Since $B_{0},B_{1}\leq B$, and $B$ is almost disjoint from every element of $\mathcal{A}$, so are $B_{0},B_{1}$ and so $B_{0},B_{1}\in\mathcal{H}^{-}(\mathcal{A})$. If $n\in T_{0}$, then $\operatorname{supp}(x_{n})\subseteq P_{n}\subseteq S_{\alpha}$, $x_{n}\in(e_{n})_{n\in S_{\alpha}}$. Therefore, $\langle B_{0}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}$. By a similar reasoning, we may also conclude that $\langle B_{1}\rangle\subseteq\langle B\rangle\cap Z_{\alpha}^{c}$.